if in an projectile motion time to flight is increase by 1% , then what will be increase in maximum height ?

Grade:11

1 Answers

Vikas TU
7 years ago
Time of Flight,
T = 2usinthetha/g
1% of time of flight is: => 0.02usinthetha/g
increased time of flight => 2.02usinthetha/g
thus new Tof flight,
T’ = 2.02usinthetha/g
Maximum height,
H/g = u^2sin^2thetha/g*g
H/g = (T’/2.02)^2
H = g*(T’/2.02)^2
and aslo find now H’ in terms of T.
then find as: (H-H’)/H * 100